# Generalizable data-driven turbulence closure modeling on unstructured grids with differentiable physics

### v2

@@ Abstract
Differentiable physical simulators are proving to be valuable tools for developing data-driven models in computational fluid dynamics (CFD). These simulators enable end-to-end training of machine learning (ML) models embedded within CFD solvers. This paradigm enables novel algorithms which combine the generalization power and low cost of physics-based simulations with the flexibility and automation of deep learning methods. In this study, we introduce a framework for embedding deep learning models within a generic finite element solver to solve the Navier-Stokes equations, specifically applying this approach to learn a subgrid scale closure with a graph neural network (GNN). We validate our method for flow over a backwards-facing step and test its performance on novel geometries, demonstrating the ability to generalize to novel geometries without sacrificing stability. Additionally, we show that our GNN-based closure model may be learned in a data-limited scenario by interpreting closure modeling as a solver-constrained optimization. Our end-to-end learning paradigm demonstrates a viable pathway for physically consistent and generalizable data-driven closure modeling across complex geometries.
keywords:
Graph neural networks , Differentiable turbulence , a-posteriori learning , Fluid dynamics , Generalization
1 Introduction
@@
Accurate computational modeling of turbulent flows represents a critical requirement for the analysis of physical systems in a variety of engineering and scientific disciplines [1]. Although the equations of fluid motion are well-understood, accurate numerical simulation of the governing Navier-Stokes equations with computational fluid dynamics (CFD) methods remains a formidable task [2] when significant turbulent behavior is observed. In such cases, solutions to the Navier-Stokes equations are characterized by chaotic and multi-scale behavior, which makes resolving the entirety of the flow field through direct numerical simulation (DNS) often impractical [3]. The number of grid points required for DNS can be shown to scale with $Re^{9/4}$, where the Reynolds number $Re$ represents the ratio of inertial to viscous forces in the flow [4]. High Reynolds number flows are often of practical interest, but simultaneously pose a challenging modeling endeavor.
@@
Developments in CFD have largely centered on reducing the computational complexity of attaining numerical solutions to computationally expensive flow configurations. Historically, major efforts have been directed towards progressing reduced-order models (ROMs) [5, 6] and turbulence closure models [7, 8, 9]. In such cases, the system complexity is decreased by modeling only a subset of the dominant modes of the system, as in the case of ROMs, or the mean flow or large scales of the flow, in the case of Reynolds-Averaged Navier-Stokes (RANS) [10] or large eddy simulation (LES) [11] respectively. Recently, more attention has been turned towards machine learning (ML) as a potential pathway for reducing the cost of CFD methods [12]. The success of deep neural networks (DNNs), in particular, has provided new strategies for interacting with flow data, for tasks such as regression, optimization, or classification. DNNs have demonstrated enormous potential to recognize and extract complex relationships in existing data, patterns which may be used for the prediction or control of new flows. This flexibility opens up numerous avenues for novel model design and the integration of ML in CFD, with a variety of approaches explored in the past few years [13]. The development of ROMs using ML was among the first efforts to tackle dynamical modeling of fluid systems with this new wave of data-driven learning [13]. Previous works have used deep learning architectures like convolutional neural networks (CNNs) [14, 15] and recurrent neural networks (RNNs) [16] to predict the spatiotemporal evolution of flow fields, or to learn nonlinear encodings of low dimensional flow features [17]. More recent advances in ML architectures have also been applied to fluid modeling, such as using generative adversarial networks (GANs) for super-resolution [18], long short-term memory networks (LSTMs) for improved dynamic prediction [19]. Notably, graph neural networks (GNNs) have demonstrated a remarkable ability to model spatiotemporal data for fluid-flow prediction tasks [20, 21, 22, 23, 24], in addition to applications from different domains including molecular dynamics [25], traffic forecasting [26], and structural mechanics [27]. Specifically, GNNs have exhibited their capability to generalize to out-of-sample datasets (for example different flow geometries) due to their ability to perfectly represent nodes and edges in unstructured meshes, and their formulation which identifies local and non-local correlations across an unstructured dataset.
@@
Many of these DNN-based modeling strategies, however, suffer from a fundamental drawback, which is that ML is ultimately interpolative and can fail catastrophically on out-of-distribution tasks. Without an interpretable framework, it can be difficult to determine if or how a model will fail when confronted with novel data. In contrast, hybrid learning approaches combine first-principles modeling built on fundamental physical assumptions with ML. By training ML models under more guaranteed physics constraints, hybrid methods can secure stronger guarantees on generalizability and stability. For example, physical symmetries and constraints can be embedded into ML models using invariant and equivariant networks [28, 29]. Turbulence closure modeling with ML represents one example of hybrid learning for CFD research. Closure models are typically partially data-driven by nature, offering a pathway to integrating ML with more fundamental first-principles knowledge. More importantly, closure models merely augment known governing laws (here the Navier Stokes equations) which must continue to be solved to machine precision albeit on coarser spatial and temporal length scales.
@@
There has been significant interest in using DNNs for turbulence modeling in the past few years [30, 31, 32]. Given their extensive use in engineering disciplines, RANS models have received considerable attention. ML has been used to identify uncertain model predictions [33], to augment existing turbulence models [34, 35, 36], and to correct discrepancies in the Reynolds stress with respect to high-fidelity simulations [37, 38]. DNNs have also been used in LES, for predicting the sub-grid scale stress that encompasses the effect of unresolved scales on the resolved scales [39, 40, 41, 42]. Common in DNN-based modeling approaches for sub-grid scale closure is the use of multi-layer perceptron architectures, which predict the closure term locally using resolved quantities. However, non-local models such as CNNs have also been explored, showing improvement over these simpler architectures [43, 44]. Turbulence modeling with ML has often taken an a-priori approach to learning, which means the closure models are trained prior to integration within a PDE solution scheme. This requires supervised data for training that must come from, typically, DNS of a ground truth flow configuration. Furthermore, performance in an a-posteriori setting is not guaranteed by good a-priori performance, as the solver introduces new artifacts, such as numerical errors and temporal effects, that cannot be captured during training [40]. Even worse, solutions from turbulence modeling can become unstable, and these issues may arise despite achieving low a-priori errors, which is due to a problem called model-data inconsistency [45]. These disadvantages have thus generally curbed the applicability of these approaches to modeling more canonical flows, where DNS is feasible. The reader is directed to [45] for a recent review on closure modeling using ML.
@@
Fortunately, the accuracy and stability in a-posteriori settings can be guaranteed by employing an a-posteriori learning approach [46], where the training procedure includes solving PDE and the solution from the PDE solver is used to train a closure model. The utilization of a differentiable PDE solver enables an end-to-end training of ML models with backpropagation, which can lead to substantially improved a-posteriori accuracy and use insight via numerical analysis [47]. However, differentiable CFD solvers are not widely available due to the nontrivial nature of changing existing CFD solvers to differentiable forms, which has hindered the investigation of this paradigm. Generally, differentiability can be realized via source-to-source transformations of existing code bases [48], implementation of a solver in an automatic differentiation (AD) framework, or exploitation of existing adjoint-capable PDE solvers [49]. Several previous efforts have leveraged these methods for improving CFD schemes. Kochkov et al. accelerated DNS computations through learned interpolations of the numerical scheme [50]. Ho et al. use the adjoint-enabled field inversion technique for improving turbulence modeling of separated flows [51]. Bezgin et al. have developed a differentiable CFD solver for two-phase flows in JAX [52]. Several other works have explored ML modeling for sub-grid closure in LES using differentiable simulations [53, 54, 55, 56]. List et al. utilize a differentiable simulator with a loss function based on turbulence physics to model closure for various evolving turbulent flow scenarios [55]. Sirignano et al. deploy a deep learning-based closure model in differentiable simulation for flows around bluff bodies at a moderate Reynolds number range and evaluates their model’s a-posteriori accuracy and stability [56]. A more recent study by Quattromini et al. utilizes GNNs to model the Reynolds stress tensor in RANS simulations for the reconstruction of the mean flow around a cylinder at low Reynolds numbers [57].
Building on these earlier studies, the main goal of this study is to introduce a GNN-based sub-grid turbulence closure for unsteady flows, which is embedded in a differentiable finite element method (FEM)-based LES solver. Especially, to explore the efficacy of a GNN-based sub-grid turbulence and differentiable solver for challenging flows, we apply our framework to a complex flow around the backwards-facing step (BFS), where near-wall turbulent flows, flow separation, and the advection of shed vortices occur simultaneously. The overall flowchart of our work is illustrated in Fig. 1. The main contributions of our study are as follows:
  • 1.
    Differentiable FEM-based LES solver: A differentiable FEM-based LES solver compatible with unstructured meshes is developed. This differentiable solver is used to generate datasets including the ground truth data and a solution from a trained model. For the baseline FEM solver, FEniCS, an open source FEM-based LES solver[58] is used. By implementing the discrete adjoint method in FEniCS, differentiability is instantiated.
  • 2.
    Generalizable sub-grid closure modeling using GNN: A sub-grid closure model using GNN is trained within the differentiable algorithm. Nodes and edges in unstructured mesh elements are represented by the nodes and edges in graphs, and non-locality of the sub-grid closure is effectively learned. A single trained model trained on the BFS case is tested on extrapolation tasks with unseen geometries including a ramp and wall-mounted cube.
  • 3.
    Training under incomplete data availability: Additional GNN-based sub-grid closure models are trained under sparse data conditions to evaluate their ability to learn closures in an inverse problem formulation. Only flowfield data downstream of the step is provided as training data, and their prediction capabilities for the entire flowfield around the step and other geometries are tested.
@@
Figure 1: The flowchart of FEM-based LES solver with differentiable algorithm for training a GNN-based closure model. $\overline{\mathbf{u}}_{i}$ and $\overline{p}_{i}$ represent velocity and pressure data from the filtered ground truth solutions, respectively. $\hat{\mathbf{u}}_{i}$ and $\hat{p}_{i}$ represent predicted velocity and pressure data. Filtered Navier-Stokes equations are solved using the IPCS algorithm [59], consisting of three steps. The GNN’s input variable $\phi$ is extracted from features at each node, and the output $\alpha_{\tau}$ includes turbulence parameters for sub-grid scale closure. The loss $\mathcal{L}$ is calculated over multiple snapshots, and gradients are backpropagated across the PDE solver to update GNN’s network parameters $\theta$ using the adjoint method. Further details about each component are explained in the Sec. 2.
@@
The remainder of this paper is organized as follows. The LES governing equations, the definitions of graph and its structures, and implementation of the adjoint method are explained in Sec. 2. Sec. 3 elaborates the data generation and training procedure. Training results and test results on unseen geometries are provided in Sec. 4, and lastly, Sec. 5 concludes this study.
2 Methods
2.1 Governing equations
@@
where $\mathbf{u}$ is the velocity vector, defined $\{u,v\}$, $p$ is the pressure, $\rho$ is the density, $Re$ is the Reynolds number, and $\mathbf{f}$ represents any external forces. The nonlinear nature of the convective term $\nabla\cdot(\mathbf{u}\otimes\mathbf{u})$ is a particularly problematic aspect of the system, resulting in chaotic solutions that are multi-scale – spanning several orders of magnitude in space and time. Accurate direct numerical simulation of the governing equations must resolve the smallest scales of the flow, which is infeasible with our current computational capabilities for all but simplified, idealized, canonical flows. The number of grid points required scales with $Re^{9/4}$ [4], which makes DNS for many high Reynolds’ number flows intractable. Fortunately, it is often sufficient in scientific and engineering applications to obtain only coarse-grained approximations of the true solution fields. The vast majority of energy in the flow is contained within the large scales, and consequently often relevant quantities, e.g. lift or drag, are influenced mainly by the scales encompassed in coarse-grained approximations.
Coarse-grained solution fields, however, cannot be recovered simply by simulating the governing equations on a coarser computational mesh. This can be shown mathematically by modeling the coarse-graining operation as a low-pass filtering operation, denoted with an overbar $\overline{\bm{\;\cdot\;}}$. Filtering, as a result of coarse-graining, leads to the introduction of additional terms in the equation, again due to the nonlinear convective term:
@@
The additional terms are representative of the “closure” problem that has plagued numerical analysis of the Navier-Stokes equations for many years, and must be modeled to effectively solve these filtered equations. The equations above represent the filtered Navier-Stokes equations used in large eddy simulation, where the unknown term is treated as an additional stress, known as the sub-grid scale (SGS) stress. We remark here, that the nature of the filtering induced by the coarse-graining is typically unknown and the rest of this manuscript will utilize filtering and coarse-graining interchangeably. Furthermore, since we are limited by no assumptions of an effective filter, we do not consider comparisons with benchmark closure modeling strategies which require approximations of low-pass filters such as dynamic Smagorinsky [60], which requires a test filter, and approximate deconvolution-based methods [61], which iteratively invert a Gaussian filter.
@@
where $C_{s}$ is a dimensionless empirical coefficient, $\Delta$ is a characteristic length scale, and $|\overline{\mathbf{S}}|=(2\overline{S}_{ij}\overline{S}_{ji})^{1/2}$ is the magnitude of the rate-of-strain tensor. The eddy viscosity assumption allows for a simple approximation of the SGS stress $\bm{\tau}^{EV}$, which is sufficient for certain flows. Further developments in SGS modeling over the years have incorporated additional nonlinear terms into the SGS stress to account for more complex flow phenomena [63, 64, 65]. Our model uses this mixed model paradigm to compute a nonlinear SGS stress as
$\hat{\bm{\tau}}=\bm{\tau}^{EV}+\bm{\tau}^{NL},$ (7)
where $\bm{\tau}^{EV}$ is computed from the Smagorinsky model and $\bm{\tau}^{NL}$ is a data-driven nonlinear contribution evaluated from a neural network.
2.2 Graph neural networks

### v3

@@ Abstract
Differentiable physical simulators are proving to be valuable tools for developing data-driven models for computational fluid dynamics (CFD). In particular, these simulators enable end-to-end training of machine learning (ML) models embedded within CFD solvers. This paradigm enables novel algorithms which combine the generalization power and low cost of physics-based simulations with the flexibility and automation of deep learning methods. In this study, we introduce a framework for embedding deep learning models within a finite element solver for incompressible Navier-Stokes equations, specifically applying this approach to learn a subgrid-scale (SGS) closure with a graph neural network (GNN). We first demonstrate the feasibility of the approach on flow over a two-dimensional backward-facing step, using it as a proof of concept to show that solver-consistent training produces stable and physically meaningful closures. Then, we extend this to a turbulent flow over a three-dimensional backward-facing step. In this setting, the GNN-based closure not only attains low prediction errors, but also recovers key turbulence statistics and preserves multiscale turbulent structures. We further demonstrate that the closure can be identified in data-limited learning scenarios as well. Overall, the proposed end-to-end learning paradigm offers a viable pathway toward physically consistent and generalizable data-driven SGS modeling on complex and unstructured domains.
[…]
I Introduction
@@
Accurate computational modeling of turbulent flows represents a critical requirement for the analysis of physical systems in a variety of engineering and scientific disciplines [1]. Although the equations of fluid motion are well-understood, accurate numerical simulation of the governing Navier-Stokes equations with computational fluid dynamics (CFD) methods remains a formidable task [2] when significant turbulent behavior is observed. In such cases, solutions to the Navier-Stokes equations are characterized by chaotic and multi-scale behavior, which makes resolving the entire flow field through direct numerical simulation (DNS) often impractical [3]. The number of grid points required for DNS can be shown to scale with $Re^{9/4}$, where the Reynolds number $Re$ represents the ratio of inertial to viscous forces in the flow [4]. High Reynolds number flows are often of practical interest, but simultaneously pose a challenging modeling endeavor.
@@
Developments in CFD have largely centered on reducing the computational complexity of attaining numerical solutions to computationally expensive flow configurations. Historically, major efforts have been directed towards progressing reduced-order models (ROMs) [5, 6] and turbulence closure models [7, 8, 9]. In such cases, the system complexity is decreased by modeling only a subset of the dominant modes of the system, as in the case of ROMs, or the mean flow or large scales of the flow, in the case of Reynolds-Averaged Navier-Stokes (RANS) [10] or large eddy simulation (LES) [11] respectively. Recently, more attention has been turned towards machine learning (ML) as a potential pathway for reducing the cost of CFD methods [12]. The success of deep neural networks (DNNs), in particular, has provided new strategies for interacting with flow data, for tasks such as regression, optimization, or classification. DNNs have demonstrated enormous potential to recognize and extract complex relationships in existing data, which may be used for the prediction or control of new flows. This flexibility opens up numerous avenues for novel model design and the integration of ML in CFD, with a variety of approaches explored in the past few years [13]. The development of ROMs using ML was among the first efforts to tackle dynamical modeling of fluid systems with this new wave of data-driven learning [13]. Previous works have used deep learning architectures like convolutional neural networks (CNNs) [14, 15] and recurrent neural networks (RNNs) [16] to predict the spatiotemporal evolution of flow fields, or to learn nonlinear encodings of low dimensional flow features [17]. More recent advances in ML architectures have also been applied to fluid modeling, such as using generative adversarial networks (GANs) for super-resolution [18], long short-term memory networks (LSTMs) for improved dynamic prediction [19]. Notably, graph neural networks (GNNs) have demonstrated a remarkable ability to model spatiotemporal data for fluid-flow prediction tasks [20, 21, 22, 23, 24, 25, 26], in addition to applications from different domains including molecular dynamics [27], traffic forecasting [28], and structural mechanics [29]. Specifically, GNNs have exhibited their capability to generalize to out-of-sample datasets (for example different flow geometries) due to their ability to perfectly represent nodes and edges in unstructured meshes, and their formulation which identifies local and non-local correlations across an unstructured dataset.
@@
Many of these DNN-based modeling strategies for direct flow-field prediction, however, suffer from a fundamental drawback, which is that ML is ultimately interpolative and can fail catastrophically on out-of-distribution tasks. Without an interpretable framework, it can be difficult to determine if or how a model will fail when confronted with novel data. In contrast, hybrid learning approaches combine first-principles modeling built on fundamental physical assumptions with ML. By training ML models under more guaranteed physics constraints, hybrid methods can secure stronger guarantees on generalizability and stability. For example, physical symmetries and constraints can be embedded into ML models using invariant and equivariant networks [30, 31]. Turbulence closure modeling with ML represents one example of hybrid learning for CFD research. Closure models are typically partially data-driven by nature, offering a pathway to integrating ML with more fundamental first-principles knowledge. More importantly, closure models merely augment known governing laws (here the Navier Stokes equations) which must continue to be solved to machine precision albeit on coarser spatial and temporal length scales.
@@
There has been significant interest in using DNNs for turbulence modeling in the past few years [32, 33, 34]. Given their extensive use in engineering disciplines, RANS models have received considerable attention. ML has been used to identify uncertain model predictions [35], to augment existing turbulence models [36, 37, 38], and to correct discrepancies in the Reynolds stress with respect to high-fidelity simulations [39, 40]. DNNs have also been used in LES, for predicting the subgrid-scale stress that encompasses the effect of unresolved scales on the resolved scales [41, 42, 43, 44]. Common in DNN-based modeling approaches for subgrid-scale closure is the use of multi-layer perceptron architectures, which predict the closure term locally using resolved quantities. However, non-local models such as CNNs, GNNs and Fourier Neural Operators (FNOs) have also been explored, showing improvement over these simpler architectures [45, 46, 47, 48, 49].
From the perspective of LES, there are two pathways to building a closure model using data-driven techniques. The first is the so-called a-priori learning approach, where the closure models are trained prior to integration within a partial differential equation (PDE) solution scheme. This requires supervised data for training that must come from, typically, DNS of a ground truth flow configuration. Furthermore, performance in an a-posteriori setting is not guaranteed by good a-priori performance, as the solver introduces new artifacts, such as numerical errors and temporal effects, that cannot be captured during training [42]. Even worse, solutions from turbulence modeling can become unstable, and these issues may arise despite achieving low a-priori errors, which is due to a problem called model-data inconsistency [50]. These disadvantages have thus generally restricted the applicability of these approaches to modeling more canonical flows, where DNS is feasible to obtain.
@@
Fortunately, the accuracy and stability in a-posteriori settings can be guaranteed by employing an a-posteriori learning approach [49], where the training procedure includes solving PDE and the solution from the PDE solver is used to train a closure model. Recent studies have demonstrated the effectiveness of deep reinforcement learning-based a-posteriori training frameworks, where the closure parameters are optimized online through direct interaction with the LES environment[51, 52]. The utilization of a differentiable PDE solver further enables an end-to-end training of ML models with backpropagation, which can lead to substantially improved a-posteriori accuracy and use insight via numerical analysis [53]. However, differentiable CFD solvers are not widely available due to the nontrivial nature of changing existing CFD solvers to differentiable forms, which has hindered the investigation of this paradigm. Generally, differentiability can be realized via source-to-source transformations of existing code bases [54], implementation of a solver in an automatic differentiation (AD) framework, or exploitation of existing adjoint-capable PDE solvers [55]. Several previous efforts have leveraged these methods for improving CFD schemes. Kochkov et al. accelerated DNS computations through learned interpolations of the numerical scheme [56]. Ho et al. use the adjoint-enabled field inversion technique for improving turbulence modeling of separated flows [57]. Bezgin et al. have developed a differentiable CFD solver for two-phase flows in JAX [58]. Fan et al. explore accelerating turbulent fluid-structure interaction (FSI) problems using a JAX-based differentiable solver framework[59]. Several other works have explored ML modeling for subgrid closure in LES using differentiable simulations [60, 61, 62, 63]. List et al. utilize a differentiable simulator with a loss function based on turbulence physics to model closure for various evolving turbulent flow scenarios [62]. Sirignano et al. deploy a deep learning-based closure model in differentiable simulation for flows around bluff bodies and evaluate their model’s a-posteriori accuracy and stability [63]. A more recent study by Quattromini et al. utilizes GNNs to model the Reynolds stress tensor in RANS simulations for the reconstruction of the mean flow around a cylinder at low Reynolds numbers [47]. While many studies on RANS modeling have employed an a-posteriori approach [64, 65, 66], turbulence modeling for LES has often taken an a-priori learning approach, where the closure models are trained prior to integration within a PDE solution scheme. The reader is directed to [50] for a recent review on closure modeling using ML. We also remark that while this review is focused on data-driven closures, similar approaches are also showing promise in the direction of wall-modeling [67, 51, 68]
Building on these earlier studies, the primary goal of this work is to introduce a subgrid-scale turbulence closure for unsteady flows on unstructured grids, embedded within a differentiable LES solver. Unlike the authors’ previous work [49], which employed the finite volume method (FVM) on structured grid simulations, the present study utilizes the finite element method (FEM). FEM offers greater flexibility in handling complex geometries and unstructured meshes, which are often necessary in real-world flow applications. Its variational formulation also naturally integrates with differentiable programming frameworks, making it a suitable choice for gradient-based optimization and learning tasks. Furthermore, a graph neural network (GNN) is adopted as the base architecture for the machine-learning-based subgrid closure model. GNNs are particularly well-suited for unstructured mesh data, as they can effectively handle non-Cartesian geometries and irregular topological structures. This makes them ideal for modeling subgrid-scale interactions across irregular grid connectivity. To evaluate the efficacy of this GNN-based closure model and differentiable solver on complex flow scenarios, we apply our framework to the canonical backward-facing step (BFS) case in both two- and three-dimensions, which features near-wall turbulence, flow separation, and the advection of shed vortices. An overview of the proposed framework is illustrated in Fig. 1. The main contributions of this study are summarized as follows:
@@
Figure 1: A schematic of the differentiable FEM-based LES solver for training a GNN-based closure model. $\overline{\mathbf{u}}_{i}$ and $\overline{p}_{i}$ represent velocity and pressure data from the filtered ground truth solutions, respectively. $\hat{\mathbf{u}}_{i}$ and $\hat{p}_{i}$ represent predicted velocity and pressure data. The LES equations are solved using the IPCS algorithm [69]. The GNN’s inputs are the node features, $\bm{V_{in}}$, and the edge features, $\bm{E_{in}}$, and the output $C_{s}$ is a spatiotemporally variable Smagorinsky coefficient for subgrid-scale closure. The loss $\mathcal{L}$ is calculated over multiple snapshots, and gradients are backpropagated across the PDE solver to update GNN parameters $\theta$ using the discrete adjoint method. Further details about each component are explained in the Sec. II.
  • Differentiable FEM-based LES solver: A differentiable FEM-based LES solver compatible with unstructured meshes is developed. This differentiable solver is used to generate datasets including the ground truth data and a solution from a trained model. We implement the solver in the open-source FEniCS/FEniCSx finite-element framework [70, 71]: the two-dimensional preliminary case is built on legacy FEniCS, while the three-dimensional LES experiments use FEniCSx. The use of FEniCSx for the three-dimensional case is motivated by its native support for distributed-memory parallelism, improved scalability on modern HPC systems, and data-oriented architecture that facilitates integration with machine learning models, which are not fully supported in the legacy version [71]. In both implementations, differentiability is instantiated by using the discrete adjoint method.
  • Generalizable subgrid closure modeling using GNN: A subgrid-scale closure model based on a GNN is trained within the proposed differentiable algorithm. In this setup, the GNN-based closure is learned in an a-posteriori manner: each training iteration solves the coarse-grained Navier-Stokes equations with the current closure, and the resulting PDE solution is then used to update the model parameters. We first apply this framework to a two-dimensional BFS as a preliminary proof of concept. We then train and evaluate the same framework on a three-dimensional BFS and further test it on an unseen geometry to assess its generalization capability.
  • Training under incomplete data availability: We also investigate whether the proposed GNN-based closure can be learned when only partial observations are available. We first demonstrate this in the two-dimensional BFS setting, where only downstream flowfield data are available. We then extend the same sparse-data setup to the three-dimensional BFS case to evaluate its performance under constrained data availability.
@@
The remainder of this paper is organized as follows. The LES governing equations, the definitions of graph and its structures, and implementation of the adjoint method are explained in Sec. II. Section III. then details the generation of the training dataset, the training procedure, the training results, the generalization experiments, and the experiments conducted under limited-data conditions. Within this section, Sec. III.1. presents the two-dimensional preliminary validation results, while Sec. III.2. provides the main three-dimensional turbulence results. Finally, Section IV. concludes the paper.
II Methods
II.1 Governing equations
@@
where $\mathbf{u}$ is the velocity vector, defined $\{u,v\}$, $p$ is the pressure, $\rho$ is the density, $Re$ is the Reynolds number, and $\mathbf{f}$ represents any external forces. The nonlinear nature of the convective term $\nabla\cdot(\mathbf{u}\otimes\mathbf{u})$ is a particularly problematic aspect of the system, resulting in chaotic solutions that are multi-scale – spanning several orders of magnitude in space and time. Accurate direct numerical simulations of the governing equations must resolve the smallest scales of the flow, which is infeasible with our current computational capabilities for all but simplified, idealized, canonical flows. The number of grid points required scales with $Re^{9/4}$ [4], which makes DNS for many high Reynolds number flows intractable. Fortunately, it is often sufficient in scientific and engineering applications to obtain only coarse-grained approximations of the true solution fields. The vast majority of energy in the flow is contained within the large scales, and consequently relevant quantities, e.g. lift or drag, are often mainly influenced by the scales encompassed in coarse-grained approximations.
Coarse-grained solution fields, however, cannot be recovered simply by simulating the governing equations on a coarser computational mesh. This can be shown mathematically by modeling the coarse-graining operation as a low-pass filtering operation, denoted with an overbar $\overline{\bm{\;\cdot\;}}$. Filtering, as a result of coarse-graining, leads to the introduction of additional terms in the equation due to the nonlinear convective term:
@@
The additional terms are representative of the “closure” problem that has plagued numerical analysis of the Navier-Stokes equations for many years, and must be modeled to effectively solve these filtered equations. The equations above represent the filtered Navier-Stokes equations used in LES, where the unknown term is treated as an additional stress, known as the subgrid-scale (SGS) stress. We remark here, that the nature of the filtering induced by the coarse-graining is typically unknown and the rest of this manuscript will utilize filtering and coarse-graining interchangeably. Furthermore, since we are limited by no assumptions of an effective filter, we do not consider comparisons with benchmark closure modeling strategies which require approximations of low-pass filters such as dynamic Smagorinsky [72], which requires a test filter, and approximate deconvolution-based methods [73], which iteratively invert a Gaussian filter.
@@
where $C_{s}$ is a dimensionless empirical coefficient, $\Delta$ is a characteristic length scale, and $|\overline{\mathbf{S}}|=(2\overline{S}_{ij}\overline{S}_{ji})^{1/2}$ is the magnitude of the rate-of-strain tensor. In our model, $C_{s}$ adaptively varies across the domain and over time as a function of both local and neighboring flow features, adjusting the SGS tensor to better match the ground truth flowfield.
II.2 Graph neural networks